1/3

Christian Morgenstern, Christian Morgenstern, Christian Morgenstern,

Christian Morgenstern,

La lupfanto- mo The Banshee (An Approach)

Der Werwolf

The Werewolf

translated by Bertilo Wennergren translated by Max Knight translated by Alexander Gross

Jen lupfantom’ de

famili’ meznokte iam ku-

ris for al tombo de la

profesor’. kaj petis pri kon-

jugaci’.

One night, a banshee

slunk away from mate and child,

and in the gloom went to a village tea-

cher’s tomb, requesting him: “Inflect

me, pray.”

Ein Werwolf eines

Nachts entwich von Weib und Kind

und sich begab an eines Dorfschul-

lehrers Grab und bat ihn: �Bitte,

beuge mich!“ A Werewolf, troubled by

his name, Left wife and brood one

night and came To a hidden graveyard to

enlist The aid of a long-dead phi-

lologist.

Do diktis tiu dok-

ta ul’ starante sur la

kructabul’ dum lup’ senmova

kiel splinto atentis vortojn de

l’ mortinto.

The village teacher

climbed up straight upon his grave stone

with its plate and to the apparition

said who meekly knelt befo-

re the dead:

Der Dorfschulmeister

stieg hinauf auf seines Blech-

schilds Messingknauf

und sprach zum Wolf,

der seine Pfoten geduldig kreuzte vor

dem Toten:

“Oh sage, wake up, please

don’t berate me,” He howled sadly, “Just con-

jugate me.” The seer arose a bit unstea-

dy Yawned twice, wheezed on-

ce, and then was ready.

De “lupfantomo”,

simpla rekta, tra “lupfintomo”,

plej perfekta, plu “lupfontomo”,

iom rara, ĝis “lupfuntomo”,

ho, koŝmara.

“The banSHEE, in the

subject’s place; the banHERS, the pos-

sessive case. The banHER, next, is

what they call objective case–and

that is all.”

�Der Werwolf“ – sprach der gute Mann,

�des Werwolfs, Geni-

tiv sodann, dem Werwolf, Dativ,

wie man’s nennt, den Werwolf, - damit

hat’s ein End!“ “Well, ‘Werewolf’ is your

plural past, While ‘Waswolf’ is singu-

larly cast: There’s ‘Amwolf’ too, the

present tense, And ‘Iswolf,’ ‘Arewolf’ in

this same sense.”

Ĉi participoj

plaĉis multe al la fantomo, kaj

rezulte ĝi petis: “Legu plu

el PIV-o pri participoj en

pasivo.”

The banshee marveled

at the cases and writhed with plea-

sure, making faces, but said: “You did not

add, so far, the plural to the singu-

lar!”

Dem Werwolf schmei-

chelten die Fälle er rollte seine Au-

genbälle. �Indessen“, bat er,

�füge doch zur Einzahl auch die

Mehrzahl noch!“ “I know that–I’m no men-

tal cripple– The future form and parti-

ciple Are what I crave,” the be-

ast replied. The scholar paused–again

he tried:

…  …  …  …

2/3

La viro ĉagreniĝis

tamen: “Jam fermis mi

per fin’ kaj amen,

ĉar netransiraj be-

stoj feblas kaj la pasiv’ ne-

niam eblas.”

The teacher, though, admitted then that this was not within his ken. “While bans are frequent”, he advised, “A she cannot be pluri-

zed.”

Der Dorfschullehrer aber musste gestehn, dass er von ihr nichts wusste. Zwar Wölfe gab’s in großer Schar, doch �Wer“ gab’s nur im Sinular.

“A ‘Will-be-wolf?’ It’s just

too long: ‘Shall-be-wolf?’ ‘Has-been-wolf?’ Utterly wrong! Such words are wounds beyond all suture– I’m sorry, but you have no

future.”

Sed lupfantomo

pri transiro ja spertas pli ol eĉ

vampiro. La lup’ revenis

larmokula al sia famili’ ulula.

The banshee, rising

clammily, wailed: “What about my family?” Then, being not a learned creature, said humbly “Thanks” and left the teacher.

Der Wolf erhob sich tränenblind – er hatte ja doch Weib und Kind! Doch da er keine Gelehrter eben, so schied er dankend und ergeben.

The Werewolf knew better–his sons still slept

At home, and homewards

now he crept, Happy, humble, without

apology For such folly of philology.

…  …  …  …

3/3

Translation of the German poem “Der Werwolf” by Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) into Esperanto by Bertilo Wennergren.

Arg-825-1653 (2013-02-14 23:29:47)

Vidu  la  re-

tejon http: // bertilow. com/ literaturo/ lupfantomo. html .

Translation of the German poem “Der Werwolf” by Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) into English by Max Knight (civila nomo: Max Kühnel, ∗1909-06-08 – †1993-08-31).

Arg-825-1660 (2013-02-13 22:36:06)

Tiu ĉi angligo troviĝas en http: // bertilow. com/ literaturo/ lupfantomo. html . Pri la tradukinto Max Knight vidu: http: // www. jbeilharz. de/ morgenstern/ morgenstern_poems. html kaj http: // www. 123people. de/ ext/ frm? ti= ersonensuche% 20telefonbuch&search_ term= m x% 20k% C3% BChnel&search_ country= DE&st= suc e% 20nach% 20personen&target_ url= aHR0cDovL d3dy5saXRlcmF0dXJoYXVzLmF0L2luZGV4LnBocD9pZD02ODc4 .

Author of this German poem is Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31).

Arg-825-1652 (2013-02-11 18:44:21)

Translation of the German poem “Der Werwolf” by Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) into English by Alexander Gross.

Arg-825-1657 (2013-02-13 18:39:45)

Tiu ĉi angligo troviĝas en http: // bertilow. com/ literaturo/ lupfantomo. html .